# A Hitchhiker’s Guide to Understanding Performances of Two-Class Classifiers

### v1

@@ Abstract
Properly understanding the performances of classifiers is essential in various scenarios. However, the literature often relies only on one or two standard scores to compare classifiers, which fails to capture the nuances of application-specific requirements, potentially leading to suboptimal classifier selection. Recently, a paper on the foundations of the theory of performance-based ranking introduced a tool, called the Tile, that organizes an infinity of ranking scores into a 2D map. Thanks to the Tile, it is now possible to evaluate and compare classifiers efficiently, displaying all possible application-specific preferences instead of having to rely on a pair of scores. In this paper, we provide a first hitchhiker’s guide for understanding the performances of two-class classifiers by presenting four scenarios, each showcasing a different user profile: a theoretical analyst, a method designer, a benchmarker, and an application developer. Particularly, we show that we can provide different interpretative flavors that are adapted to the user’s needs by mapping different values on the Tile. As an illustration, we leverage the newly introduced Tile tool and the different flavors to rank and analyze the performances of $74$ state-of-the-art semantic segmentation models in two-class classification through the eyes of the four user profiles. Through these user profiles, we demonstrate that the Tile effectively captures the behavior of classifiers in a single visualization, while accommodating an infinite number of ranking scores.  [^1] This paper is the third of a trilogy. In a nutshell, paper A [44] presents an axiomatic framework and an infinite family of scores for ranking classifiers. In paper B [45], we particularize this framework to binary classification and present the Tile that organizes these scores (among which $PPV$, $TPR$, $TNR$, $F_{1}$, and $A$) in a single plot. Finally, this paper (paper C [26]) provides a guide to using the Tile according to four practical scenarios. For that, we present different Tile flavors that are applied to a real case.
@@ 1 Introduction
As humans, performance and ranking are widespread in all aspects of our lives. For instance, in school, teachers evaluate tests and homework using a score which reflects the performance of students. In some disciplines such as calculus, evaluation is straightforward as there are only two possible cases: either the answer is correct or wrong. The score can then be calculated as the ratio of correct answers to the total number of questions, which is straightforward. Likewise, in most team sports, team A beats team B if they score more points. Even for reviewing papers, area chairs use scores provided by the reviewers to assess if a paper should be accepted to or rejected from the conference [38].
@@
In this paper, we propose a step-by-step hitchhiker’s guide to help compare, analyze, and rank two-class classifiers. Throughout this guide, we study four scenarios, answering specific requests that four common user profiles may have, and provide visual tools adapted to their needs, as illustrated in Fig. 1. To do so, we rely on the infinite and parametric family of scores satisfying an axiomatic definition of the performance-based rankings introduced in paper A [44] and the tool called the Tile introduced in paper B [45]. More precisely, we leverage the newly introduced Tile tool to present different flavors for visualizing various elements useful for understanding the performances of two-class classifiers under different angles and considering an infinite number of scores. As such, we propose several ways to construct, use, and interpret the Tile in the four scenarios, offering practical guidance to different user profiles with varied objectives, such as theoretically analyzing scores, developing new methods, benchmarking challenges, or designing applications. Finally, we show that the Tile is a versatile tool that can serve a wide range of people in the field of computer vision through an illustration that analyzes and ranks $74$ state-of-the-art semantic segmentation models trained on $4$ datasets.
@@ 2 Related Work
In general terms, Japkowicz and Shah [33] decompose performance evaluation and analysis into four parts: (1) the analysis of performance measures, (2) the estimation of errors, (3) the study of statistical testing, and (4) understanding of the experimental power of datasets. In machine learning, the importance of these parts is related to the application. For example, when datasets are too small, the amount of uncertainty in the measured performances should be scrutinized with tools that examine their statistical significance [18]. However, statistical significance is less of an issue in the field of computer vision, as it commonly disposes of enough samples to evaluate a classifier. For the analysis of performance measures, which is the focus of our guide, two strategies have emerged.
The first strategy consists in providing a series of performance scores that highlight the benefits and trade-offs in designing a classifier. Tharwat [52] analyzed a series of scores for classification tasks that can be used alone or in combination. Hereafter, we discuss some of these scores. For example, the weighted harmonic mean between the recall $TPR$ and the precision $PPV$, denoted by $F_{\beta}$, or simply by $F_{1}$ when weights are equal, is advocated by many authors. A score similar to $F_{1}$ is Jaccard’s coefficient. Yet, it is known that Jaccard’s coefficient and the $F_{1}$ score lead to the same ranking of classifiers, although Jaccard is probabilistic, unlike $F_{1}$ [23, 47]. The unweighted arithmetic mean of $TPR$ and $TNR$ is the balanced accuracy. It is considered as the most suitable score for evaluating imbalanced problems, thanks to its independence regarding the priors [3, 7, 25, 36]. However, the problem with the first strategy is that even though we can use some scores for analyzing the performances of classifiers and ranking them, it is not clear which ones to use.
The second strategy makes use of so-called evaluation spaces, such as the Receiver Operating Characteristic (ROC) [41] or Precision-Recall (PR) spaces, that combine two scores. The ROC space is a plot of sensitivity vs. one minus specificity as one varies a cutoff on a continuous predictor used to decide [27]. In [21, 22], Fawcett explained how to use ROC graphs and avoid interpretation pitfalls. For instance, the area under the ROC curve (AUC-ROC) is commonly used as a performance indicator, as it has statistical significance [2]. Although several authors raised concerns about ROC-derived scores in decision-making [6, 10], recent works such as the one on contingency spaces [1] still build upon ROC graphs. More recently, the PR space has become the de facto replacement for the ROC space in the presence of imbalance [32], ignoring the fact that there is a bijection between these two spaces [17]. Moreover, the PR space suffers from the presence of an unachievable region [4], which is often overlooked in practice. The primary drawback of ROC- or PR-based curves analyzes is that, according to Menon et al. [40], the AUC-ROC and AUC-PR apply not to a classifier but to a scoring function, and that a scoring function yields a family of classifiers obtained for different thresholds. In other words, these curves consider a parametric family of classifiers rather than a specific classifier, and they are therefore inadequate to select classifiers when it comes to comparing unique instances from different families.
@@
As an example of these drawbacks, we show the performance of $74$ two-class classifiers in the ROC space in Fig. 2. As can be seen, it is extremely challenging to uniquely determine which classifier is the best or the worst. Moreover, it is even harder to rank all $74$ classifiers based only on this ROC space. For instance, relying on a single of the two presented scores such as the True Positive Rate ($TPR$) may result in overfitting on that score, disregarding other important nuances of the performance.
While these strategies provide a good insight from multiple perspectives, they only offer a scattered interpretation due to the inherent partial redundancy between scores, and ultimately only offer an incomplete look on the performances and ranking. In a recent attempt to formalize the notion of ranking, Nguyen et al. [42] proposed to impose three properties for ranking: (1) reliability, (2) meaningfulness (evaluated by humans), and (3) mathematical consistency. In paper A [44], we present an alternative formal axiomatic definition of performance-based rankings, grounded in order theory and anchored within a probabilistic framework. The axioms ensure the stability of rankings, i.e., that if multiple entities, in our case two-class classifiers, are ranked, adding or removing an entity does not affect the relative order of the previously present entities. Additionally, it introduces ranking scores that satisfy these axioms, parameterized by a random variable $I$, called importance, which allows for consideration of application-specific preferences. Paper B [45] proposes a spatial organization of the ranking scores in a 2D square map, called the Tile, for the particular case of two-class classification. It also studies properties of the Tile from a theoretical perspective. In this work, we show how both the axiomatic framework and the Tile can be used in practice and propose a hitchhiker’s guide to use the Tile in various ways through four scenarios and user profiles.
@@ 3 Hitchhiker’s Guide
In this section, we present our hitchhiker’s guide to understanding performances of two-class classifiers. We begin by providing the recommended basic theoretical knowledge for using the guide, including a description of the terminology and notations, the general Tile tool, and the illustrative context used throughout. Next, we detail four scenarios, each aligned with the needs of a specific user profile. For each scenario, we provide the context, highlight the relevant flavors of the Tile, and explain how to interpret the results.
@@ 3.1 Recommended basic theoretical knowledge
Terminology and notations. To be consistent with paper A [44] and paper B [45], we use the same terminology and notations. Hence, an entity $\epsilon$ is a two-class classifier and the set of all entities of interest is denoted by $\mathbb{E}$. Performances, denoted by $P$, are probability measures, and the performance $P_{\epsilon}$ of an entity $\epsilon$ is the evaluation of this entity. Scores are functions associating a real value to performances, that is, $X:\mathrm{dom}(X)\rightarrow\mathbb{R}:P\mapsto X(P)$, with the domain of the score, $\mathrm{dom}(X)$, included in the set of all probability measures, $\mathbb{P}_{(\Omega,\Sigma)}$, on the measurable space $(\Omega,\Sigma)$, where $\Omega$ is the sample space (i.e., the set of outcomes) and $\Sigma$ is the event space (i.e., a $\sigma$-algebra on $\Omega$).
Construction and interpretation of the Tile. As explained in paper B [45], the Tile for two-class classification is obtained by organizing ranking scores on a 2D square map, where the two axes, $a$ and $b$, respectively represents the importance $I$ given to true positive ($tp$) compared to true negative ($tn$), and false negative ($fn$) compared to false positive ($fp$):
@@
with $P(\{tp\})$ (resp. $P(\{tn\})$, $P(\{fp\})$, $P(\{fn\})$, and $P(\{tn\})$) corresponding to the probability of event $\{tp\}$ (resp. $\{fp\}$, $\{fn\}$, and $\{tn\}$).
The Tile is then defined as follows:
@@
with $\mathbb{X}_{(\Omega,\Sigma)}$ denoting all possible scores on $(\Omega,\Sigma)$ and $I(tn)=1-a$, $I(fp)=1-b$, $I(fn)=b$, $I(tp)=a$ ,
@@
Figure 3: Tile with canonical ranking scores for two-class classification. Each point of the Tile corresponds to a ranking score which can be computed using Eq. 3, with the two axes, $a$ and $b$, defined in Eqs. 1 and 2, respectively. The Tile therefore contains an infinity of ranking scores, including the popular Accuracy ($A$), Negative Predictive Value ($NPV$), True Positive Rate ($TPR$), Positive Predictive Value ($PPV$), True Negative Rate ($TNR$), and $F_{1}$ scores.
@@
Table 1: List of the models (columns) sorted in chronological order based on the date of publication (from 2015 to 2023) from the MMSegmentation toolbox [15] trained on one or multiple datasets (rows). We respectively have $31$ models trained on the ♠ Cityscapes [16] dataset, $27$ on the ♥ ADE20K [70] dataset, $12$ on the ♠ Pascal VOC 2012 [19] dataset, and $4$ on the ♣ COCO-Stuff 164k [8] dataset, totaling $74$ models.
@@
Description of our illustrative context. We illustrate each scenario with a semantic segmentation example. For evaluating the performance of entities, we use the BDD100K dataset [62] as a testing set and define a two-class problem by taking the union of the first $11$ semantic classes of the dataset as the negative class, corresponding to background objects (e.g., road, sidewalk, or building), and the union of the $8$ remaining classes as the positive class, corresponding to foreground, potentially moving, objects (e.g., person, rider, or car). This type of binary split has been commonly used to improve motion detection algorithms [5, 43].
For the choice of entities, we compare and rank $74$ state-of-the-art semantic segmentation models, trained on different learning sets and available in the MMSegmentation toolbox [15]. As shown in Tab. 1, we gather respectively $31$ models trained on ♠ Cityscapes [16], $27$ on ♥ ADE20K [70], $12$ on ♠ Pascal VOC 2012 [19], and $4$ on ♣ COCO-Stuff 164k [8]. The classes used as learning sets are also grouped into a positive and a negative class, following the same logic as for BDD100K. Since the testing set is different from all learning sets, we are therefore evaluating the generalization capability of semantic segmentation models in a two-class setting. An overview of the results is presented through different flavors of the Tile that illustrate the scenarios. For more details, we provide a complete report on all $74$ models in the supplementary material.
@@ 3.2 Four Scenarios of our Hitchhiker’s Guide
In this section, we provide a step-by-step guide for using the Tile to rank two-class classifiers (i.e., entities), by illustrating its various potential uses through four scenarios. Each scenario corresponds to one user profile among (1) the theoretical analyst, who is interested in understanding the theoretical relationship between different scores typically used for evaluating or ranking methods, (2) the method designer, who needs to analyze the performances of his/her new method and compare it to others, (3) the benchmarker, who organizes challenges for the scientific community and has to rank participating methods, and finally (4) the application developer, who should be able to select the most appropriate method for his/her application. In the final scenario, we also discuss various strategies for selecting an entity based on the Tile.
Scenario 1: The Theoretical Analyst
@@
Scenario 2: The Method Designer
@@
In practice, several methods can be used to compute the Value Tile, depending on the availability of data. (1) Direct computation: If the values of $P(\{fp\})$, $P(\{fn\})$, $P(\{tp\})$, and $P(\{tn\})$ are known, the first method consists in computing the values of the ranking scores in each point of the Tile using Eq. 3. (2) Interpolation: Knowing the values of the scores $TPR$, $TNR$, $NPV$, and $PPV$ at the four corners of the Tile, the values of the remaining scores can be determined by averaging, using the $f$-mean: in particular, vertically, the harmonic mean $f:x\mapsto x^{-1}$ and, horizontally, an f-mean defined as $f:x\mapsto(1-x)^{-1}$. As one can see, the non-linearities differ between the vertical and horizontal axes. (3) Equation system resolution: The third method involves first, determining the values of $P(\{fp\})$, $P(\{fn\})$, $P(\{tp\})$, and $P(\{tn\})$ by solving a system of $4$ equations with $4$ unknowns, then, computing the values of the ranking scores using Eq. 3. Specifically, if the values of three ranking scores from the Value Tile are known, or if two ranking scores and the priors (for fixed priors) with $\pi_{-}=P(\{tn\})+P(\{fp\})$ and $\pi_{+}=P(\{fn\})+P(\{tp\})$, then, knowing that
$P(\{fp\})+P(\{fn\})+P(\{tp\})+P(\{tn\})=1$ (10)
makes the system complete. Let us note that all three methods require discretizing the Tile. This can be implemented with a grid size parameter (set to $2001$ in our code), which defines linearly spaced values for the two axes, $a$ and $b$.
@@
Scenario 3: The Benchmarker
@@
Ranking Tile. The Ranking Tile maps, for a given entity $\epsilon$, the rank of that entity across the tile. For a given set of entities $\mathbb{E}$, rankings are based on an ordering of the performances, defined in paper A [44], induced by the ranking scores $R_{I}$, denoted by $\lesssim_{R_{I}}$. The Ranking Tile is defined as:
@@
Therefore, the Ranking Tile can be obtained by ordering the performances for each ranking score, i.e., ordering in each point $(a,b)$ the values of the ranking score for the Value Tile of each entity of the set. This means that the Ranking Tile is discretized similarly to the Value Tile.
@@
Scenario 4: The Application Developer
@@
The Entity Tile thus shows all entities that are at a given rank and is constructed similarly to the Ranking Tile. In the special case of $r=1$, the Entity Tile shows the best entities among $\mathbb{E}$, while for $r=|\mathbb{E}|$, it shows the worst entities among $\mathbb{E}$. As illustrated in Fig. 8 for the semantic segmentation models, the first rank is shared by $4$ entities, namely Mask2Former and SETR, both trained on ♠ Cityscapes, and DeepLabV3+ and ISANet, both trained on ♠ Pascal VOC 2012. These $4$ entities also appear in the second and third ranks in other areas of the ranking tiles, i.e., for other canonical ranking scores. Interestingly, the properties ensure that entities ranked first are the ones on the upper dashed broken line of Fig. 2 in the ROC space, and the ones ranked last are on the lower dashed broken line  [^2] Further details on the link between the Tile and ROC are provided in paper B [45]..

### v3

@@ Abstract
Properly understanding the performances of classifiers is essential in various scenarios. However, the literature often relies only on one or two standard scores to compare classifiers, which fails to capture the nuances of application-specific requirements. The Tile is a recently introduced visualization tool organizing an infinity of ranking scores into a 2D map. Thanks to the Tile, it is now possible to compare classifiers efficiently, displaying all possible application-specific preferences instead of having to rely on a pair of scores. This hitchhiker’s guide to understanding the performances of two-class classifiers presents four scenarios showcasing different user profiles: a theoretical analyst, a method designer, a benchmarker, and an application developer. We introduce several interpretative flavors adapted to the user’s needs by mapping different values on the Tile. We illustrate this guide by ranking and analyzing the performances of $74$ state-of-the-art semantic segmentation models through the perspective of the four scenarios. Through these user profiles, we demonstrate that the Tile effectively captures the behavior of classifiers in a single visualization, while accommodating an infinite number of ranking scores. Code for mapping the different Tile flavors is available in supplementary material.
@@ 1 Introduction
As humans, performance and ranking are widespread in all aspects of our lives. For instance, at school, teachers evaluate tests using a score which reflects the performance of students. In some disciplines such as calculus, evaluation is straightforward as there are only two possible cases: either the answer is correct or wrong. The score can then be calculated as the ratio of correct answers to the total number of questions. Likewise, in most team sports, team A beats team B if they score more points. Even for reviewing papers, area chairs use scores provided by the reviewers to assess if a paper should be accepted or rejected [30].
@@
In this paper, we propose a step-by-step hitchhiker’s guide to help compare, analyze, and rank two-class classifiers. Throughout this guide, we study four scenarios, answering specific requests that four common user profiles may have, and provide visual tools adapted to their needs, as illustrated in Fig. 1. To do so, we rely on the infinite and parametric family of scores satisfying an axiomatic definition of the performance-based rankings introduced in [37] and the tool called the Tile introduced in [36]. More precisely, we leverage the newly introduced Tile tool to visualize various elements useful for understanding the performances of two-class classifiers from different perspective and considering an infinite number of scores. As such, we propose several ways to construct, use, and interpret different Tile flavors in the four scenarios, offering practical guidance to different user profiles with varied objectives, such as theoretically analyzing scores, developing new methods, benchmarking challenges, or designing applications.
We provide Python code that takes a list of two-class classification performances and generates a report (see supplementary material) from which most of the figures in this guide are issued. This report is intended for the theoretical analyst [Uncaptioned image], the method designer [Uncaptioned image], the benchmarker [Uncaptioned image], and the application developer [Uncaptioned image]. It contains a recap of the compared performances (tables and plots), all flavors described in this guide (Correlation Tiles, Value Tiles, Baseline Value Tile, State-of-the-Art Value Tile, Ranking Tiles, and Entity Tile) as well as others (No-Skill Tile, Relative-Skill Tile), and gives several recommendations for the models to select (see scenario 4). Besides generating a report, our code also provides implementations of algorithms introduced in [36].
@@ 2 Related Work
For the analysis of performance measures, which is the focus of our guide, two strategies have emerged.
The first strategy consists in providing a series of performance scores that highlight the benefits and trade-offs in designing a classifier. Tharwat [45] analyzed a series of scores for classification tasks that can be used alone or in combination. For example, the weighted harmonic mean between the recall $TPR$ and the precision $PPV$, denoted by $F_{\beta}$, or simply by $F_{1}$ when weights are equal, is advocated by many authors. A score similar to $F_{1}$ is Jaccard’s coefficient. Yet, it is known that Jaccard’s coefficient and the $F_{1}$ score lead to the same ranking of classifiers, although Jaccard is probabilistic, unlike $F_{1}$ [20, 39]. However, the problem with the first strategy is that even though we can use some scores for analyzing the performances of classifiers and ranking them, it is not clear which ones to use.
The second strategy makes use of so-called evaluation spaces, such as the Receiver Operating Characteristic (ROC) [33] or Precision-Recall (PR) spaces, that combine two scores. In [18, 19], Fawcett explained how to use ROC graphs and avoid interpretation pitfalls. For instance, the area under the ROC curve (AUC-ROC) is commonly used as a performance indicator, as it has statistical significance [2]. Although several authors raised concerns about ROC-derived scores in decision-making [4, 8], recent works such as the one on contingency spaces [1] still build upon ROC graphs. More recently, the PR space has become the de facto replacement for the ROC space in the presence of imbalance [26], ignoring the fact that there is a bijection between these two spaces [15]. Moreover, the PR space suffers from the presence of an unachievable region [3], which is often overlooked in practice. The primary drawback of ROC- or PR-based curves analyzes is that, according to Menon et al. [32], the AUC-ROC and AUC-PR apply not to a classifier but to a scoring function, and that a scoring function yields a family of classifiers obtained for different thresholds. In other words, these curves consider a parametric family of classifiers rather than a specific classifier, and they are therefore inadequate to select classifiers when it comes to comparing unique instances from different families.
@@
As an example of these drawbacks, we show the performance of $74$ two-class classifiers in the ROC space in Fig. 2. As can be seen, it is extremely challenging to uniquely determine which classifier is the best or the worst. Moreover, it is even harder to rank all $74$ classifiers based only on this ROC space.
While these strategies provide a good insight from multiple perspectives, they only offer a scattered interpretation due to the inherent partial redundancy between scores, and ultimately only offer an incomplete look on the performances and ranking. In a recent attempt to formalize the notion of ranking, Nguyen et al. [34] proposed to impose three properties for ranking: (1) reliability, (2) meaningfulness (evaluated by humans), and (3) mathematical consistency. Piérard et al. [37] presents an alternative formal axiomatic definition of performance-based rankings, grounded in order theory and anchored within a probabilistic framework. The axioms ensure the stability of rankings, i.e., that if multiple entities, in our case two-class classifiers, are ranked, adding or removing an entity does not affect the relative order of the previously present entities. Additionally, it introduces ranking scores that satisfy these axioms, parameterized by a random variable $I$, called importance, which allows for consideration of application-specific preferences. Piérard et al. [36] proposes a spatial organization of the ranking scores in a 2D square map, called the Tile, for the particular case of two-class classification. It also studies properties of the Tile from a theoretical perspective. In this work, we show how both the axiomatic framework and the Tile can be used in practice and propose a hitchhiker’s guide to use the Tile in various ways through four scenarios and user profiles.
@@ 3 Hitchhiker’s Guide
This section presents our hitchhiker’s guide to understanding performances of two-class classifiers. We begin by providing the recommended basic theoretical knowledge for using the guide, including a description of the terminology and notations, the general Tile tool, and the illustrative context used throughout. Next, we detail four scenarios, each aligned with the needs of a specific user profile. For each scenario, we highlight the relevant flavors of the Tile, and explain how to interpret the results.
@@ 3.1 Recommended Basic Theoretical Knowledge
Terminology and notations. To be consistent with [37] and [36], we use the same terminology and notations. Hence, an entity $\epsilon$ is a two-class classifier and the set of all entities of interest is denoted by $\mathbb{E}$. Performances, denoted by $P$, are probability measures, and the performance $P_{\epsilon}$ of an entity $\epsilon$ is the evaluation of this entity. Scores are functions associating a real value to performances, that is, $X:\mathrm{dom}(X)\rightarrow\mathbb{R}:P\mapsto X(P)$, with the domain of the score, $\mathrm{dom}(X)$, included in the set of all probability measures, $\mathbb{P}_{(\Omega,\Sigma)}$, on the measurable space $(\Omega,\Sigma)$, where $\Omega$ is the sample space (i.e., the set of outcomes) and $\Sigma$ is the event space (i.e., a $\sigma$-algebra on $\Omega$).
Construction and interpretation of the Tile. As explained in [36], the Tile for two-class classification is obtained by organizing ranking scores on a 2D square map, where the two axes, $a$ and $b$, respectively represents the importance $I$ given to true positive ($tp$) compared to true negative ($tn$), and false negative ($fn$) compared to false positive ($fp$):
@@
with $P(\{tn\})$ (resp. $P(\{fp\})$, $P(\{fn\})$, and $P(\{tp\})$) corresponding to the probability of the event $\{tn\}$ (resp. $\{fp\}$, $\{fn\}$, and $\{tp\}$). The Tile is then defined as follows:
@@
with $\mathbb{X}_{(\Omega,\Sigma)}$ denoting all possible scores on $(\Omega,\Sigma)$ and $I(tn)=1-a$, $I(fp)=1-b$, $I(fn)=b$, $I(tp)=a$.
@@
Figure 3: Tile with ranking scores for two-class classification. Each point of the Tile corresponds to a ranking score $R_{I}$ which can be computed using Eq. 3. The Tile therefore contains an infinity of ranking scores, including the popular Accuracy ($A$), Negative Predictive Value ($NPV$), True Positive Rate ($TPR$), Positive Predictive Value ($PPV$), True Negative Rate ($TNR$), and $F_{1}$ scores. The location of a ranking score $R_{I}$ in the Tile is defined by the values of $a$ and $b$, which indicate the importance given to true positives ($tp$), true negatives ($tn$), false negatives ($fn$), and false positives ($fp$). In all following figures, the Tile flavors adopt the same layout as the one on the left hand side. For ease of reading, the names of the scores at the four corners are systematically written.
@@
Table 1: List of the models (columns) sorted w.r.t. the date of publication (from 2015 to 2023) from the MMSegmentation toolbox [13] trained on one or multiple datasets (rows). We respectively have $31$ models trained on the ♠ Cityscapes [14] dataset, $27$ on the ♥ ADE20K [63] dataset, $12$ on the ♠ Pascal VOC 2012 [16] dataset, and $4$ on the ♣ COCO-Stuff 164k [6] dataset, totaling $74$ models.
@@
Description of our illustration. In the following, we illustrate all our Tile flavors on a same use case related to computer vision for autonomous driving. We demonstrate that the applicability of the Tile and its flavors is far from being limited to two-class classification problems. Specifically, we compare and rank $74$ state-of-the-art semantic segmentation models from the MMSegmentation toolbox [13]. What makes this use case challenging is that the sets of predicted classes differ from one model to another. This is because the models have been trained on several datasets with heterogeneous labels. As shown in Tab. 1, these include $31$ models trained on ♠ Cityscapes [14], $27$ on ♥ ADE20K [63], $12$ on ♠ Pascal VOC 2012 [16], and $4$ on ♣ COCO-Stuff 164k [6]. Motivated by the preponderant role played by the foreground-background differentiation in autonomous driving, we decided to group the semantic labels into a negative class, corresponding to background objects (e.g., road, sidewalk, or building), and a positive class, corresponding to foreground, potentially moving, objects (e.g., person, rider, or car). The grouping has been performed for both the learning sets and BDD100K [55], that we selected as testing set (see supplementary material for a detailed explanation on how it was performed). By doing so, the performances of the $74$ models become directly comparable.
@@ 3.2 Four Scenarios of our Hitchhiker’s Guide
This section provides a step-by-step guide for using the Tile to rank two-class classifiers (i.e., entities), by illustrating its various potential uses through four scenarios. Each scenario corresponds to one user profile among (1) the theoretical analyst [Uncaptioned image], who is interested in understanding the theoretical relationship between different scores typically used for evaluating or ranking methods, (2) the method designer [Uncaptioned image], who needs to analyze the performances of his/her new method and compare it to others, (3) the benchmarker [Uncaptioned image], who organizes challenges for the scientific community and has to rank participating methods, and finally (4) the application developer [Uncaptioned image], who should be able to select the most appropriate method for his/her application. In the final scenario, we also discuss various strategies for selecting an entity based on the Tiles.
Scenario 1: The Theoretical Analyst [Uncaptioned image]
@@
Scenario 2: The Method Designer [Uncaptioned image]
@@
In practice, several methods can be used to compute the Value Tile, which are discussed in supplementary material.
@@
The Value Tile shows a negative, zero, or positive horizontal gradient $\frac{\partial R_{I}}{\partial a}$ when, respectively, $P(\{tp\})$ is less than, equal to or greater than $P(\{tn\})$. Similarly, the vertical gradient $\frac{\partial R_{I}}{\partial b}$ is negative, zero, or positive when, respectively, $P(\{fn\})$ is less than, equal to, or greater than $P(\{fp\})$. As we have $P(\{tp\})<P(\{tn\})$ for all the models in our particular illustrative case, it is enough to look at the direction of the iso-value lines in order to analyze the bias (in the sense of Byrt et al. [5]): a model is biased towards the negative class (see, e.g., ISANet ♠ and DeepLabV3+ ♠ in Fig. 5), unbiased, or biased towards the positive class (see, e.g., SETR ♠ and Mask2Former ♠ in Fig. 5) when, respectively, the iso-value lines are tilted to the left, vertical, or tilted to the right.
@@
Scenario 3: The Benchmarker [Uncaptioned image]
@@
Ranking Tile. The Ranking Tile maps, for a given entity $\epsilon$, the rank of that entity across the tile. Following [37], for a given set of entities $\mathbb{E}$, rankings are based on an ordering $\lesssim_{R_{I}}$ of the performances induced by the ranking scores $R_{I}$. The Ranking Tile is defined as:
@@
Therefore, the Ranking Tile can be obtained by ordering the performances for each ranking score, i.e., ordering in each point $(a,b)$ the values of the ranking score for the Value Tile of each entity of the set.
@@
Scenario 4: The Application Developer [Uncaptioned image]
@@
The Entity Tile thus shows all entities that are at a given rank and is constructed similarly to the Ranking Tile. In the special case of $r=1$, the Entity Tile shows the best entities among $\mathbb{E}$, while for $r=|\mathbb{E}|$, it shows the worst entities among $\mathbb{E}$. As illustrated in Fig. 8 for the semantic segmentation models, the first rank is shared by $4$ entities, namely Mask2Former and SETR, both trained on ♠ Cityscapes, and DeepLabV3+ and ISANet, both trained on ♠ Pascal VOC 2012. These $4$ entities also appear in the second and third ranks in other areas of the Ranking Tiles, i.e., for other canonical ranking scores. Interestingly, the entities ranked first are the ones on the upper dashed broken line of Fig. 2 in the ROC space, and the ones ranked last are on the lower dashed broken line. This is a consequence of properties studied in [36].
@@
4 Discussion
Bands. Unlike the Value Tiles, all Correlation, Ranking and Entity Tiles illustrating this guide exhibit roughly horizontal bands. They can be explained as follows. All performances such that $P(\{tn\})\propto P(\{tp\})$ are perfectly rank-correlated with all ranking scores on any given horizontal line of the Tile. In our illustration, the ratio $\frac{P(\{tp\})}{P(\{tn\})}$ is roughly constant ($\mu=0.12$, $\sigma=0.01$), which explains the roughly horizontal bands. Similarly, all performances such that $P(\{fp\})\propto P(\{fn\})$ are perfectly rank-correlated with all ranking scores on any given vertical line of the Tile. This covers the case, not illustrated in this guide, in which all compared performances are unbiased, in the sense of Byrt et al. [5], i.e. $P(Y)=P(\hat{Y}$).
Multi-class classification extension. The universal theoretical foundations for performance-based ranking [37], on top of which the Tile was build [36], are indeed applicable to multi-class problems. With $n$ classes, specifying the relative importance given to the samples $\omega$ such that $\hat{Y}(\omega)=Y(\omega)$ (as done on the horizontal axis of the Tile) requires $n-1$ dimensions, and specifying the relative importance given to the samples $\omega$ such that $\hat{Y}(\omega)\neq Y(\omega)$ (as done on the vertical axis of the Tile) requires $n^{2}-n-1$ dimensions. A "ranking space" similar to the Tile, but for $n$ classes, would therefore require $n^{2}-2$ dimensions. This can be handled algorithmically, but it cannot be easily visualized. A typical alternative is to convert multi-class problems into two-class problems by micro- or macro-averaging [43]. While micro- and macro-averaged Value Tiles can be constructed and used to derive Entity Tiles, this approach is not necessarily advisable as it raises serious concerns. On the one hand, all micro-averaged scores are only functions of the multi-class accuracy, which leads to a loss of information, and, on the other hand, macro-averaging comes with interpretability issues [35]. We hope that this will motivate further studies on the topic.
Choice of a positive class. Once the labels have been grouped into two classes, there remains to choose which is the positive one, in order to comply with the classical convention. However, the opposite choice leads to identical Tiles, up to a central symmetry [36]. In particular, the list of entities (classifiers) identified as being the best (ranked first) does not depend on which class is chosen as positive.